A floor 8m 20cm by 6m is to be covered with a tiles of size 20cm by15cm. Find the number of tiles required and cost of the tiles come at ₹240 for 15 tiles.
Answer by Alan3354(69443)

A floor 8m 20cm by 6m is to be covered with a tiles of size 20cm by15cm. Find the number of tiles required and cost of the tiles come at ₹240 for 15 tiles.
------------
Area of the floor is 820*600 = 492000 sq cm
Area of a tile is 10*15 = 300 sq cm
If you're allowed to cut the tiles to fit if needed:
492000/300 = 1640 tiles